OpenStudy (anonymous):

What polynomial has roots of -5, -4, and 1?

OpenStudy (anonymous):

answer choices: x^3 - 8x^2 - 11x + 20 x^3 - x^2 + 22x + 40 x^3 + x^2 - 22x - 40 x^3 + 8x^2 + 11x - 20

OpenStudy (anonymous):

Plug each of them into each equation. Whichever one is 0 for those 3 numbers is your answer.

OpenStudy (the_fizicx99):

(x + 5)(x + 4)(x - 1) (x^2 + 9x + 20)(x - 1) = x^3 - x^2 + 9x^2 - 9x + 20x - 20 x^3 + 8x^2 + 11x - 20

OpenStudy (the_fizicx99):

I guess working backwards ~

OpenStudy (cwrw238):

it has to be the one which ends in -20

OpenStudy (the_fizicx99):

Or you could do that ^ ;-;

OpenStudy (anonymous):

ok thank you guys! i know how to do it now yay(:

OpenStudy (anonymous):

so it would have to be D

OpenStudy (cwrw238):

yes as the roots are -5, -4 and 1 the factors must be (x+5)(x+4)(x-1) and 5 * 4 * -1 = -20

OpenStudy (anonymous):

ok got it thanks again!

OpenStudy (cwrw238):

i deleted the other post because it was wrong...

OpenStudy (cwrw238):

yw